# KAN we improve on HEP classification tasks? Kolmogorov-Arnold Networks applied to an LHC physics example

### v1

@@ 5 Conclusions
As long as the KANs have a hidden layer with multiple nodes, we observe very similar performance to MLPs. Numerically, our best KAN is even slightly better than the best MLP, but this difference is very small and most likely practically irrelevant when considering systematic uncertainties in a physics analysis. These findings seem to contradict Ref. [11], where a clear improvement over the performance of MLPs was found in several examples. In these examples, loss values were reached that were orders of magnitude lower than those of MLPs, for example in fitting $f(x,y)=x\cdot y$. However, those examples are of much lower dimensionality and complexity than our classification task. We suppose that the examples in Ref. [11] are especially well-suited for learning the Kolmogorov-Arnold representation of the underlying functional relationship. However, our dataset includes features with a stronger variability in shape. The representations that have to be learned to solve our classification task may hence not be particularly suited for the KAN architecture.
@@
In terms of interpretability, we find that small KANs indeed offer advantages. For a one-layer KAN, we observe that the learned activation functions resemble the log-likelihood ratios of the input features. In addition, the $L_{1}$-norm of the activation functions offers a straightforward interpretation of the importance of different input features. Somewhat larger KANs may still offer an illustrative visualization of the activation functions and their $L_{1}$-norms. However, for KANs of greater depth or with wider layers, interpretability seems challenging.
Because of their better interpretability than MLPs, we conclude that KANs are a promising alternative for classification tasks in HEP if the performance of small KANs is sufficient or if moderate performance losses are acceptable in favor of interpretability. We believe that more research on the application of KANs in HEP tasks is necessary. In particular, prospects for the interpretability of larger KANs should be explored.

### v2

@@ 5 Conclusions
As long as the KANs have a hidden layer with multiple nodes, we observe very similar performance to MLPs. Numerically, our best KAN is even slightly better than the best MLP, but this difference is very small and most likely practically irrelevant when considering systematic uncertainties in a physics analysis. These findings seem to contradict Ref. [17], where a clear improvement over the performance of MLPs was found in several examples. In these examples, loss values were reached that were orders of magnitude lower than those of MLPs, for example, in fitting $f(x,y)=x\cdot y$. However, those examples are of much lower dimensionality and complexity than our classification task. We suppose that the examples in Ref. [17] are especially well-suited for learning the Kolmogorov-Arnold representation of the underlying functional relationship. However, our dataset includes features with a stronger variability in shape. The representations that have to be learned to solve our classification task may hence not be particularly suited for the KAN architecture.
@@
In terms of interpretability, we find that small KANs indeed offer advantages. For a one-layer KAN, we observe that the learned activation functions resemble the univariate log-likelihood ratios of the input features which these functions act on. In addition, the $L_{1}$-norms of the activation functions offer a straightforward interpretation of the importance of different input features for such KANs with only one layer. Somewhat larger KANs may still offer an illustrative visualization of the activation functions, which we regard as an intrinsic interpretability advantage of KANs. However, for KANs of greater depth or with wider layers, interpretability seems challenging.
Because of their better interpretability compared to MLPs, we conclude that KANs are a promising alternative for classification tasks in HEP when the performance of small KANs is sufficient or when moderate performance losses are acceptable in favor of interpretability. We believe that more research on applying KANs in HEP tasks is necessary. This study primarily focuses on the performance and parameter efficiency of KANs compared to MLPs, as well as the interpretability of small KANs. The interpretability of deeper KANs should be explored in future studies to address this limitation of our current work. Mechanistic interpretability techniques may provide systematic ways to understand the learned representations in these more complex models, and exploring symbolic regression to approximate the learned activation functions could open new avenues for applying KANs as a tool in HEP. In addition, assessing KAN performance and interpretability in HEP regression tasks may further broaden their applicability beyond classification.